OpenStudy (anonymous):

2x+(5/6)=3x-(5/6) solve for x

OpenStudy (saifoo.khan):

Raheen to the Rescue.!

OpenStudy (anonymous):

x= 5/3

OpenStudy (anonymous):

omg u guys are insanely good at this

OpenStudy (anonymous):

combine like terms 2x+(5/6)=3x-(5/6) 10/6=x x=5/3

OpenStudy (anonymous):

such equation you may grasp it directly

OpenStudy (diyadiya):

\[2x + \frac{5}{6} = 3x - \frac{5}{6}\] \[\frac{12 x+ 5}{6} = \frac{18x-5}{6}\] 12x + 5 = 18x - 5 18x - 12x = 10 6x = 10 \[x=\frac{5}{3}\] : D